OpenStudy (anonymous):

Solve. –2(6 – 20n) = 20(2n – 3)

OpenStudy (anonymous):

–2(6 – 20n) = 20(2n – 3) so multiply 6 - 20n with -2 and 2n - 3 with 20 which is -12 + 40n = 40n - 60 hope it helped

OpenStudy (anonymous):

@Audry,hemp ... please donot use such language...! its a study site!

OpenStudy (anonymous):

LEARN LANGUAGE THAT ALSO GOOD ONE OKAY BABY